Prove that maximum value of a^2*b^3*c^4 subject to a+b+c=18 is 4^2*6^3*8^4.


1).....without using  AM,GM equality thing...



2).......using  AM,GM equality thing...


PLS EXPLAIN!!!!!!

Grade:9

1 Answers

Aman Bansal
12 years ago

Dear Shivam,

Arithmetic mean is always greater than or equal to geometric mean

A.M. >= G.M.

Lets say a,b,c are 3 number such that a+b+c=m

Now you have to find maximum value of (a^2) x (b^3) x(c^4)


Since it is given that a+b+c=m

=> a/2 +a/2 +b/3 +b/3 +b/3 +c/4 +c/4 +c/4 +c/4 = m

=> (a/2 +a/2 +b/3 +b/3 +b/3 +c/4 +c/4 +c/4 +c/4)/9 = m/9

We know that AM >= GM

Max value of GM= AM, which is attained when all the numbers are equal

Max of [ (a/2)^2 x (b/3)^3 x(c/4)^4] ^1/9= (a/2 +a/2 +b/3 +b/3 +b/3 +c/4 +c/4 +c/4 +c/4)/9 = m/9

And for maxima, a/2=b/3=c/4

And max (a/2)^2 x (b/3)^3 x(c/4)^4 = (m/9)^9

max of (a)^2 x (b)^3 x(c)^4= (2^2) (3^3)(4^4) (m/9)^9

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert